Currently the cost of platinum is $1,544.70 per ounce. If a pure palladium coin displaces 0.591 mL in a graduated cylinder, how much is this coin worth? (Given: 1 oz = 28.35 g, 1 mL = 1 cm3, Density
1
Expert's answer
2020-07-30T07:08:33-0400

Find mass of a coin:

"m=V\times d = 0.591 cm^3 \times(\frac{24.45 g}{1 cm^3})= 14.45 g"


Covert gram to oz:

"14.45 g \times (\frac{1 oz}{28.35 g})=0.5097 oz"


Find cost of a coin:

"0.5097 g \times (\frac{1544.70 }{1 oz})= 787.3" $